Solve the Schrödinger equation for a particle in a box with sides at 𝑥 = −𝐿 and 𝑥 = 𝐿. Determine the eigenvalues and the normalized eigenfunctions. 


1
Expert's answer
2021-09-15T11:39:29-0400

The Schrödinger equation for wave function of the system is as follows

"-\hbar^2/2m\psi''(x)+V(x)\psi(x)=E\psi(x)"

For particle inside a box "V(x)=0", so

"-\hbar^2/2m\psi''(x)=E\psi(x)"

Solution

"\psi(x)=A\cos(kx)+B\sin(kx), \quad k=\sqrt{2mE}/\hbar"

The boundary conditions give

"\psi(L)=A\cos(kL)+B\sin(kL)=0,\\
\psi(-L)=A\cos(kL)-B\sin(kL)=0."

So

"A=0,\:\quad kL=\pi n,\quad E_n=\frac{\hbar^2\pi^2n^2}{2mL^2}"

"\psi_n(x)=B\sin(\pi nx/L)"

Normalization condition gives

"\int_{-L}^L\psi^2(x)dx=1"

So

"B^2\int_{-L}^L\sin^2(\pi n x/L)dx=1,\quad B=1/\sqrt{L}"

Finally

"\psi_n(x)=1/\sqrt{L}\sin(\pi nx/L)"